# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 5.1. Characterizations of topological freeness
If $C_{0}(X)$ is maximal abelian in some exotic groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$, then $\mathcal{G}$ is topologically free. If $G$ is topologically free and Hausdorff, then $C_{0}(X)$ is maximal abelian in every reduced groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
Assume that $\mathcal{G}$ is not topologically free, so that there is a non-empty open bisection $V\subseteq G\setminus X$ for which $r|_{V}=d|_{V}$. Thus every $b\in C_{c}(V,\mathcal{L})\subseteq F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ commutes with $C_{0}(X)$ for any algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ that contain the spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$ (so for instance an exotic Banach algebra). Hence $C_{0}(X)$ is not maximal abelian in $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
It is known that in the untwisted case topological freeness is equivalent to the generalized intersection property for essential groupoid $C^{*}$-algebras, see [KM21, Theorem 7.29]. This generalizes classical results for crossed products by Kawamura–Tomiyama [KT90] and Archbold–Spielberg [AS93]. A corresponding result for $L^{P}$-operator algebra crossed products was proved in [BK24, Theorem 5.9]. We now generalize all these results to essential groupoid Banach algebras associated to $P\subseteq[1,\infty]$. To this end, we use the following representation, which in the $\ell^{2}$-context is called the orbit representation [KM21], augmentation representation [BCFS14], or trivial representation [Ren97]. As in [BK24] we prefer to call it a $C_{0}(X)$-trivial representation.
@@
For any $p\in[1,\infty]$, we have a representation $\Lambda^{\operatorname{tr}}_{p}:F_{p}({\mathcal{G}})\to B(\ell^{p}(X))$ where
$\Lambda^{\operatorname{tr}}_{p}(f)\xi(x):=\sum_{d(\gamma)=x}f(\gamma)\xi(r(\gamma)),\qquad f\in\mathfrak{C}_{c}(\mathcal{G}),\ \xi\in\ell_{p}(X).$
@@
Consider the canonical action $h:S\to\operatorname{PAut}(X)$ of any wide inverse semigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$, cf. [BKM23, Example 2.17]. One readily sees that
@@
for $a\in C_{0}(X)$, $\xi\in\ell^{p}(X)$, $x\in X$, $U\in S$, defines a covariant representation $(\pi^{\operatorname{tr}},v^{\operatorname{tr}})$ of $h$ on $\ell^{p}(X)$ in the sense of [BKM23, Definition 3.18]. The integrated representation $\pi^{\operatorname{tr}}\rtimes v^{\operatorname{tr}}:F({\mathcal{G}})\to B(\ell^{p}(X))$, see [BKM23, Proposition 4.21], satisfies the formula in the assertion and it descends to the desired representation $\Lambda^{\operatorname{tr}}_{p}:F_{p}({\mathcal{G}})\to B(\ell^{p}(X))$. ∎
Recall that for any non-empty $P\subseteq[1,\infty]$ we denoted by $\Lambda_{P}:F^{P}(\mathcal{G},\mathcal{L})\to F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ the canonical regular representation, see Definition 3.14. We denote by $j_{P}^{\operatorname{e}}:F^{P}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ the canonical contractive linear map, which is identity on every $C_{0}(U,\mathcal{L})$, $U\in\operatorname{Bis}(G)$, cf. Remark 4.14. Following [BK24], we also define the $C_{0}(X)$-trivial representation $\Lambda_{P}^{\operatorname{tr}}$ of $F^{P}(\mathcal{G})$ as an extension of the $\ell^{\infty}$-direct sum $\oplus_{p\in P}\Lambda_{p}^{\operatorname{tr}}$ of representations from Lemma 5.12.
@@
Let $\mathcal{G}$ be an étale groupoid. For any non-empty $P\subseteq[1,\infty]$ the following conditions are equivalent:
@@
    $\ker(\Lambda_{P}^{\operatorname{tr}})\subseteq\ker(j_{P}^{\operatorname{e}})$ where $\Lambda_{P}^{\operatorname{tr}}$ is the $C_{0}(X)$-trivial representation of $F^{P}(\mathcal{G})$;
@@
(1) implies (2) by Theorem 5.10, and (2) implies (3) because $\ker(\Lambda_{P}^{\operatorname{tr}})$ is an ideal in $F^{P}(\mathcal{G})$ satisfying $\ker(\Lambda_{P}^{\operatorname{tr}})\cap C_{0}(X)=\{0\}$. To show that (3) implies (1) assume that $\mathcal{G}$ is not topologically free. Thus there is a non-empty open bisection $U\subseteq\mathcal{G}\setminus X$ with $r|_{U}=d|_{U}$. Since $U\cap X=0$, we get $\operatorname{\mathbb{E}}_{X}(f)=0$ for all $f\in C_{0}(U)$. Choose any non-zero $f\in C_{c}(U)$ and define $f_{0}\in C_{c}(d(U))\subseteq C_{0}(X)$ by $f_{0}(d(\gamma)):=f(\gamma)$ for $\gamma\in U$. Since $r|_{U}=d|_{U}$, both $\Lambda^{\operatorname{tr}}_{P}(f)$ and $\Lambda^{\operatorname{tr}}_{P}(f_{0})$ act on each subspace $\ell^{p}(X)$, $p\in P$, by pointwise multiplication with the same function $f_{0}$. Hence $f-f_{0}\in\ker\Lambda^{\operatorname{tr}}_{P}$. For the canonical essential expectation $\operatorname{\mathbb{E}}_{X}^{P}:F^{P}(\mathcal{G})\to\mathcal{D}(X)$ we have $\operatorname{\mathbb{E}}_{X}^{P}(f-f_{0})=-f_{0}\neq 0$. Hence $f-f_{0}\notin\ker(j_{P}^{\operatorname{e}})$, cf. Proposition 4.12. Thus (2) fails.
This proves that (1)(3) are equivalent. These conditions are independent of the choice of $P$ because (1) is. Assuming $\mathfrak{M}_{0}(\mathcal{G})=\{0\}$ we have $\ker(j_{P}^{\operatorname{e}})=\ker j_{P}=\ker(\Lambda_{P})$, by Remark 4.13 and Proposition 3.15. For a non-empty $Q\subseteq\{1,\infty\}$ we have $\ker(\Lambda_{Q})=\{0\}$, by Remark 3.16. Hence condition (4) is equivalent to (2) for $P=Q$, because $F^{Q}(\mathcal{G})=F^{Q}_{\operatorname{r}}(\mathcal{G})=F^{Q}_{\operatorname{e}}(\mathcal{G})$. ∎
@@
When $\mathcal{G}$ is a countable union of open bisections, then $\ker(j_{P}^{\operatorname{e}})$ coincides with the kernel of the essential representation $\widetilde{\Lambda}_{P}^{\operatorname{e}}:F^{P}(\mathcal{G},\mathcal{L})\to\widetilde{F}^{P}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ given by the direct sum $\ell^{\infty}$-direct sum $\oplus_{p\in P}\widetilde{\Lambda}_{p}$ of representations from Proposition 4.16, see also Corollary 4.19.
@@
Assume in addition that $\mathcal{G}$ is second countable and amenable, then for any non-empty set $P\subseteq[1,\infty]$ the above are further equivalent to:
@@
If $\mathfrak{M}_{0}(\mathcal{G})=\{0\}$, then all reduced algebras are essential, see Remark 4.13. Hence (1) implies (2) by Theorem 5.10. Implication (2$\Rightarrow$(3) is obvious because the algebras $F^{1}(\mathcal{G})$, $F^{\infty}(\mathcal{G})$, $F_{I}(\mathcal{G})$ are reduced. If $\mathcal{G}$ is second countable and amenable, then $F^{P}(\mathcal{G})=F^{P}_{\operatorname{r}}(\mathcal{G})$, by [GL17, Theorem 6.19], and hence these algebras are reduced. Thus also (2$\Rightarrow$(4) in the amenable case. Since simplicity implies both the intersection property and minimality, see Lemmas 5.2 and 5.4, we get that either of (3) or (4) implies (1) by Theorem 5.13. ∎
In the twisted case we have the following characterization of topological freeness that generalizes [BK24, Theorem 5.13] (formulated for group actions) and [KM22, Proposition 6.4] in the case of a line bundle.
@@
(1) implies (2) by Theorem 5.10, because topological freeness of $\mathcal{G}$ passes to an open subgroupoid $\mathcal{H}\subseteq\mathcal{G}$ containing $X$. Implication (2$\Rightarrow$(3) is obvious. To prove that (3) implies (1) assume that (1) fails. Then there is a non-empty bisection $U\subseteq\mathcal{G}\setminus\overline{X}$ with $r|_{U}=d|_{U}$. Then the union $\mathcal{H}:=X\cup\bigcup_{k\in\mathbb{Z}}U^{k}$ is an open subgroupoid of $\mathcal{G}$ generated by $X$ and $U$. The open set $U^{0}:=d(U)=r(U)$ is $\mathcal{H}$-invariant. In particular, the groupoid $\mathcal{H}_{0}=\bigcup_{k\in\mathbb{Z}}U^{k}$ is an ideal in $\mathcal{H}$, and so denoting by $B$ the closure of the image of $\mathfrak{C}_{c}(\mathcal{H}_{0},\mathcal{L}|_{\mathcal{H}_{0}})$ in $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ we see that $B$ is an ideal. In fact, $B$ is an ideal in $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ generated by $C_{0}(U^{0})=B\cap C_{0}(X)$.
@@
Thus to show that (3) fails (which finishes the proof) it is enough to show that $C_{0}(U^{0})\subseteq B$ does not have the intersection property. We have the following two cases:
1) If $U^{n}\cap U^{0}=U^{n}\cap X=\emptyset$ for all $n>0$, then $\{U^{n}\}_{n\in\mathbb{Z}}$ are pairwise disjoint and $\mathcal{H}_{0}\cong U^{0}\times\mathbb{Z}$ is the transformation groupoid for the trivial $\mathbb{Z}$-action (since $U$ is a bisection with $r|_{U}=d|_{U}$ we have $U^{n}:=\{\gamma^{n}:\gamma\in U\}$).
@@
Thus we may assume that $\mathcal{H}_{0}$ is the transformation groupoid for the trivial $G$-action where $G=\mathbb{Z}$ or $G=\mathbb{Z}_{n}$. Then $\mathcal{H}$ is a transformation groupoid for a partial $G$-action generated by $\textrm{id}|_{U_{0}}$, cf. Remark 7.2 below. In particular, $\mathcal{H}$ is Hausdorff and $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})=F^{P}_{\operatorname{r}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ is a reduced groupoid algebra. This implies that $B\subseteq F^{P}_{\operatorname{r}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ can be naturally identified with the reduced algebra $F^{P}_{\operatorname{r}}(\mathcal{H}_{0},\mathcal{L}|_{\mathcal{H}_{0}})$. Thus $B$ can be identified with a reduced $L^{P}$-operator algebra crossed product studied in [BK24]. Since the group $G$ is amenable, the trivial action is amenable and $B$ is in fact also a full $L^{P}$-operator algebra crossed product by [BK24, Corollary 4.14]. Since the $G$-action is trivial, by passing to a subgroup of $G$ if necessary, we get that $C_{0}(U^{0})\subseteq B$ does not have the intersection property by [BK24, Corollary 5.13]. (Formally, in [BK24] it was assumed that $\mathbb{F}=\mathbb{C}$ but a quick inspection of the proofs of the invoked results show that they remain valid for $\mathbb{F}=\mathbb{R}$, also due to our definition of the universal algebras in the real case.) ∎

### v3

@@ 5.2. Characterizations of topological freeness
If $C_{0}(X)$ is maximal abelian in some exotic groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$, then $\mathcal{G}$ is topologically free. If $\mathcal{G}$ is topologically free and Hausdorff, then $C_{0}(X)$ is maximal abelian in every reduced groupoid Banach algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
Assume that $\mathcal{G}$ is not topologically free, so that there is a non-empty open bisection $V\subseteq\mathcal{G}\setminus X$ for which $r|_{V}=d|_{V}$. Thus every $b\in C_{c}(V,\mathcal{L})\subseteq F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ commutes with $C_{0}(X)$ for any algebra $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ that contain the spaces $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$ (so for instance an exotic Banach algebra). Hence $C_{0}(X)$ is not maximal abelian in $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$.
@@
It is known that, in the untwisted case, topological freeness is equivalent to the generalized intersection property for essential groupoid $C^{*}$-algebras, see [KM21, Theorem 7.29]. This generalizes classical results for crossed products by Kawamura–Tomiyama [KT90] and Archbold–Spielberg [AS93]. A corresponding result for $L^{P}$-operator algebra crossed products was proved in [BK24, Theorem 5.9]. We now generalize all these results to essential groupoid Banach algebras associated to $P\subseteq[1,\infty]$. To this end, we use the following representation, which in the $\ell^{2}$-context is called the orbit representation [KM21], augmentation representation [BCFS14], or trivial representation [Ren97]. As in [BK24], we prefer to call it a $C_{0}(X)$-trivial representation.
@@
For any $p\in[1,\infty]$, we have a representation $\Lambda^{\operatorname{tr}}_{p}:F^{p}({\mathcal{G}})\to B(\ell^{p}(X))$ where
$\Lambda^{\operatorname{tr}}_{p}(f)\xi(x):=\sum_{d(\gamma)=x}f(\gamma)\xi(r(\gamma)),\qquad f\in\mathfrak{C}_{c}(\mathcal{G}),\ \xi\in\ell^{p}(X).$
@@
Consider the canonical action $h:S\to\operatorname{PAut}(X)$ of any wide inverse semigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$, cf. [BKM25, Example 2.19]. One readily sees that
@@
for $a\in C_{0}(X)$, $\xi\in\ell^{p}(X)$, $x\in X$, $U\in S$, defines a covariant representation $(\pi^{\operatorname{tr}},v^{\operatorname{tr}})$ of $h$ on $\ell^{p}(X)$ in the sense of [BKM25, Definition 3.17]. By [BKM25, Proposition 4.23 and Theorem 5.5], $(\pi^{\operatorname{tr}},v^{\operatorname{tr}})$ integrates to a $\|\cdot\|_{I}$-contractive homomorphism $\pi^{\operatorname{tr}}\rtimes v^{\operatorname{tr}}:\mathfrak{C}_{c}(\mathcal{G})\to B(\ell^{p}(X))$. Hence it extends to the desired representation $\Lambda^{\operatorname{tr}}_{p}:F^{p}({\mathcal{G}})\to B(\ell^{p}(X))$, see Example 2.6. ∎
Recall that for any non-empty $P\subseteq[1,\infty]$ we denoted by $\Lambda_{P}:F^{P}(\mathcal{G},\mathcal{L})\to F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ the canonical regular representation, see Definition 3.14. We denote by $j_{P}^{\operatorname{e}}:F^{P}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ the canonical contractive linear map, which is the identity on every $C_{0}(U,\mathcal{L})$, $U\in\operatorname{Bis}(G)$, cf. Remark 4.14. Following [BK24], we also define the $C_{0}(X)$-trivial representation $\Lambda_{P}^{\operatorname{tr}}$ of $F^{P}(\mathcal{G})$ as an extension of the $\ell^{\infty}$-direct sum $\oplus_{p\in P}\Lambda_{p}^{\operatorname{tr}}$ of representations from Lemma 5.12.
@@
Let $\mathcal{G}$ be an étale groupoid. For any non-empty $P\subseteq[1,\infty]$, the following conditions are equivalent:
@@
    $\ker(\Lambda_{P}^{\operatorname{tr}})\subseteq\ker(j_{P}^{\operatorname{e}})$ where $\Lambda_{P}^{\operatorname{tr}}$ is the $C_{0}(X)$-trivial representation of $F^{P}(\mathcal{G})$.
@@
(1) implies (2) by Theorem 5.10, and (2) implies (3) because $\ker(\Lambda_{P}^{\operatorname{tr}})$ is an ideal in $F^{P}(\mathcal{G})$ satisfying $\ker(\Lambda_{P}^{\operatorname{tr}})\cap C_{0}(X)=\{0\}$. To show that (3) implies (1), assume that $\mathcal{G}$ is not topologically free. Thus there is a non-empty open bisection $U\subseteq\mathcal{G}\setminus X$ with $r|_{U}=d|_{U}$. Since $U\cap X=\emptyset$, we get $\operatorname{\mathbb{E}}_{X}(f)=0$ for all $f\in C_{0}(U)$. Choose any non-zero $f\in C_{c}(U)$ and define $f_{0}\in C_{c}(d(U))\subseteq C_{0}(X)$ by $f_{0}(d(\gamma)):=f(\gamma)$ for $\gamma\in U$. Since $r|_{U}=d|_{U}$, both $\Lambda^{\operatorname{tr}}_{P}(f)$ and $\Lambda^{\operatorname{tr}}_{P}(f_{0})$ act on each subspace $\ell^{p}(X)$, $p\in P$, by pointwise multiplication with the same function $f_{0}$. Hence $f-f_{0}\in\ker\Lambda^{\operatorname{tr}}_{P}$. For the canonical essential expectation $\operatorname{\mathbb{E}}_{X}^{P}:F^{P}(\mathcal{G})\to\mathcal{D}(X)$ we have $\operatorname{\mathbb{E}}_{X}^{P}(f-f_{0})=-f_{0}\neq 0$. Hence $f-f_{0}\notin\ker(j_{P}^{\operatorname{e}})$, cf. Proposition 4.12. Thus (2) fails.
This proves that (1)(3) are equivalent. These conditions are independent of the choice of $P$ because (1) is. Assuming $\mathfrak{M}_{0}(\mathcal{G})=\{0\}$, we have $\ker(j_{P}^{\operatorname{e}})=\ker j_{P}=\ker(\Lambda_{P})$, by Remark 4.13 and Proposition 3.15. For a non-empty $Q\subseteq\{1,\infty\}$, we have $\ker(\Lambda_{Q})=\{0\}$, by Remark 3.16. Hence condition (4) is equivalent to (2) for $P=Q$, because $F^{Q}(\mathcal{G})=F^{Q}_{\operatorname{r}}(\mathcal{G})=F^{Q}_{\operatorname{e}}(\mathcal{G})$. ∎
@@
When $\mathcal{G}$ is a countable union of open bisections, then $\ker(j_{P}^{\operatorname{e}})$ coincides with the kernel of the essential representation $\widetilde{\Lambda}_{P}^{\operatorname{e}}:F^{P}(\mathcal{G},\mathcal{L})\to\widetilde{F}^{P}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ given by the $\ell^{\infty}$-direct sum $\oplus_{p\in P}\widetilde{\Lambda}_{p}$ of representations from Proposition 4.16, see also Corollary 4.19.
@@
Assume in addition that $\mathcal{G}$ is second countable and amenable, then for any non-empty set $P\subseteq[1,\infty]$, the above are further equivalent to:
@@
[…]
In the twisted case, we have the following characterization of topological freeness that generalizes [BK24, Theorem 5.13], which is formulated for group actions, and [KM22, Proposition 6.4] in the case of a line bundle.
@@
(1) implies (2) by Theorem 5.10, because topological freeness of $\mathcal{G}$ passes to an open subgroupoid $\mathcal{H}\subseteq\mathcal{G}$ containing $X$. It is clear that (2) implies (3). To prove that (3) implies (1), assume that (1) fails. Then there is a non-empty bisection $U\subseteq\mathcal{G}\setminus\overline{X}$ with $r|_{U}=d|_{U}$. Then the union $\mathcal{H}:=X\cup\bigcup_{k\in\mathbb{Z}}U^{k}$ is an open subgroupoid of $\mathcal{G}$ generated by $X$ and $U$. The open set $U^{0}:=d(U)=r(U)$ is $\mathcal{H}$-invariant. In particular, the groupoid $\mathcal{H}_{0}=\bigcup_{k\in\mathbb{Z}}U^{k}$ is an ideal in $\mathcal{H}$ (that is, $\mathcal{H}\mathcal{H}_{0}\subseteq\mathcal{H}_{0}$ and $\mathcal{H}_{0}\mathcal{H}\subseteq\mathcal{H}_{0}$), and so if $B$ denotes the closure of the image of $\mathfrak{C}_{c}(\mathcal{H}_{0},\mathcal{L}|_{\mathcal{H}_{0}})$ in $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$, we see that $B$ is an ideal. In fact, $B$ is an ideal in $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ generated by $C_{0}(U^{0})=B\cap C_{0}(X)$.
@@
Thus to show that (3) fails, which finishes the proof, it is enough to show that $C_{0}(U^{0})\subseteq B$ does not have the intersection property. We have the following two cases:
1) If $U^{n}\cap U^{0}=U^{n}\cap X=\emptyset$ for all $n>0$, then $\{U^{n}\}_{n\in\mathbb{Z}}$ are pairwise disjoint and $\mathcal{H}_{0}\cong U^{0}\times\mathbb{Z}$ is the transformation groupoid for the trivial $\mathbb{Z}$-action (since $U$ is a bisection with $r|_{U}=d|_{U}$, we have $U^{n}:=\{\gamma^{n}:\gamma\in U\}$).
@@
Thus we may assume that $\mathcal{H}_{0}$ is the transformation groupoid for the trivial $G$-action where $G=\mathbb{Z}$ or $G=\mathbb{Z}_{n}$. Then $\mathcal{H}$ is a transformation groupoid for a partial $G$-action generated by $\textrm{id}_{U^{0}}$, cf. Example 7.2 below. In particular, $\mathcal{H}$ is Hausdorff and $F^{P}_{\operatorname{e}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})=F^{P}_{\operatorname{r}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ is a reduced groupoid algebra. This implies that $B\subseteq F^{P}_{\operatorname{r}}(\mathcal{H},\mathcal{L}|_{\mathcal{H}})$ can be naturally identified with the reduced algebra $F^{P}_{\operatorname{r}}(\mathcal{H}_{0},\mathcal{L}|_{\mathcal{H}_{0}})$. Thus $B$ can be identified with a reduced $L^{P}$-operator algebra crossed product studied in [BK24]. Since the group $G$ is amenable, the trivial action is amenable and $B$ is in fact also a full $L^{P}$-operator algebra crossed product by [BK24, Corollary 4.14]. Since the $G$-action is trivial, by passing to a subgroup of $G$ if necessary, we get that $C_{0}(U^{0})\subseteq B$ does not have the intersection property by [BK24, Corollary 5.13]. Formally, in [BK24] it was assumed that $\mathbb{F}=\mathbb{C}$, but a quick inspection of the proofs of the invoked results show that they remain valid for $\mathbb{F}=\mathbb{R}$, also due to our definition of the universal algebras in the real case. ∎